1 Name three phase learning format and explain each.

2.How can a learner learn to wtrite two_digit in a way that is connected to the base 10 meaning of ones and tens.

3. Difine number and numeral and provide examples for each.

4. Draw Dienes to show how to find the splution to:

a) 56+48

b) 90_36

5. Test the following numbers for divisibility by 6,9 and 11( do not divide or factorise.)

a) 6 184 046

b) 21 589 434


1
Expert's answer
2021-05-27T06:51:41-0400

(1)Before phase: This is the launch phase which ensures students understand problem they need to solve.

During phase: Here students explore . Students evaluate and analyze problem with little supervision.

After phase: This phase allows critical thinking where students argue their answers and justify the solution.


(2) A learner learn to write two-digit in a way that is connected to the base 10 meaning of ones and tens by first grouping the set of numbers into tens to aid in counting. Secondly recognize how many tens and ones are there then write. For example when saying writing "forty six" you would write "four tens and six ones"


(3)Number is an idea and numerical is how to write the idea. For example number can be one while numerical can be 1 or one.


(4)a) 56+48


b) 90_36



(5)a)

Sum of the digits "6 +1+8+4+ +0+4+6=29" which is not divisible by 3 thus, 6184046 is not divisible by 6.

The sum of the digits is not divisible by 9 so "6184046" is not divisible by 9.

Sum of odd digits: "6+8+0+6=20"

Sum of even digits:"1+4+4=9"

Since sum of odd digits does not equal to sum of even digits, "6184046" is not divisible by 11.

b)

Sum of digits:"2+1+5+8+9+4+3+4=36" is divisible by 3 thus, "21589434" is divisible by 6.

The sum of the digits is divisible by 9 so "21589434" is divisible by 9.

Sum of odd digits: "2+5+9+3=19"

Sum of even digits: "1+8+4+4=17"

Since sum of odd digits does not equal to sum of even digits, "21589434" is not divisible by 11